(2 points) An experiment consists of flipping a coin, rolling a 16 sided die, and spinning a roulette wheel. What is the probability that the coin comes up heads and the die comes up less than 4 and the roulette wheel comes up with a number greater than 15? Ansie(#
Answer #1

Case 1:

P(Head in coin) = 1/2

(Two sides: Head & Tail. So, P(H) = 1/2)

Case 2:

P(less than 4 in 16 sided die) = 3/16

(Less than 4 are: 1, 2 and 3. So, P(Less than 4)= 3/16)

Case 3:

P(Number greater than 15 in roulette wheel) = 22/37

(Total numbers = 37. Numbers greater than 15 = 22. So, P(Number greater than 15) = 22/37)

So,

P(Head in coin AND Less than 4 in die AND Greater than 15 in Roulette wheel)= (1/2)X (3/16) X (22/37) = 66/1184=0.0557

So,

Answer is:

0.0557